OpenStudy (anonymous):

a. Determine whether the ratios 10/20 and 30/60 can form a proportion by finding a common multiplier. b. Show that the ratios in part (a) are equal by writing them in simplest form. Please explain how you got the answers to these questions.

OpenStudy (jdoe0001):

what's a ratio?

OpenStudy (anonymous):

10/20 can be changed to have a 60 on the bottom: 10*3 = 30 and 20*3 = 60. 30/60 =30/60. So they do form a proportion. b. 10/20 in its simplest form: what factor do 10 and 20 have in common? How about 10? So 10/10 = 1 and 20/10 = 2. So the final fraction is 1/2. 30/60 also needs to equal 1/2. So what do 30 and 60 have in common? 30! So 30/30 = 1 and 60/30 = 2. 1/2 is the final fraction too! 10/20 = 1/2 and 30/60 = 1/2!

OpenStudy (anonymous):

@Samantha_Rose You copied this from someone else..............